# Context-aware knowledge graph framework for traffic speed forecasting using graph neural network

### v1

@@ III-C2 Relation-dependent integration for context representations
The other component of the proposed relation-dependent integration strategy is attribute augmentation to enhance the semantic richness of each fact. This augmentation employs attribute facts associated with each fact in $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$, such as $hasPoi[Type]InBuffer[Dist]$ for $\mathbb{K}_{S}(poi)$ and $hasLand[Type]InBuffer[Dist]$ for $\mathbb{K}_{S}(land)$, infusing additional attribute information into the embeddings. For simplicity, we denote the attribute value for relations as $x$ and for entities as $y$. Consequently, the attribute-augmented embeddings for relations are $x_{1}\textbf{r}_{1},...,x_{l}\textbf{r}_{l}$, and for entities, it is $y\textbf{e}$. To ensure the consistency and comparability across attributes, we apply max-min normalization to these attribute values separately, resulting in normalized values $x^{\prime}$ and $y^{\prime}$. Integrating these normalized attributes into the relation-dependent embedding yields the enriched context-aware embedding for $e$:
@@ III-D Context-aware traffic forecast using multi-head self-attention
In this study, we propose a novel method to apply MHSA from two views, i.e., context-based and sequence-based views, enabling a dual perspective analysis. Initially, the dual-view MHSA discerns the influence of various context features in $\mathbb{K}$, prioritizing those critical for forecasting from the context-based feature view. Subsequently, the sequence-based view employs the output from the context-based view to highlight the historical sequences vital for temporal prediction, emphasizing time slots that are most predictive for future traffic conditions. In addition, forward masking is employed in the sequence-based view to ensure temporal coherence by restricting attention to past and present data. This operation prevents the model from accessing future information, which is essential to avoid lookahead bias. This dual application of MHSA augments the capacity to leverage both contextual features and historical sequences.
Upon deriving the fused context-aware representations from the dual-view MHSA, they are subsequently integrated into a graph convolutional network to facilitate traffic speed forecasting, i.e., CKG-GNN. Specifically, we leverage traffic speed data alongside the context-aware representations from the preceding $a$ time steps to forecast traffic speed for the forthcoming $b$ time steps. In this study, the diffusion convolution recurrent neural network (DCRNN) is employed as the backbone graph convolution model for traffic speed prediction, as detailed by [29]. It is important to note that while DCRNN serves as the backbone of the CKG-GNN in this context, the methodology is designed to be model-agnostic, allowing for the substitution or integration of alternative time-series prediction models. This flexibility underscores the objective to showcase the efficacy of a CKG-based framework in enhancing traffic forecasting accuracy, thereby demonstrating the potential benefits of incorporating context awareness into predictive models in the transportation domain.
@@ IV-A Datasets and processing
We sourced the traffic speed dataset from HERE technologies, focusing on Singapore’s Core Central Region (CCR), i.e., the downtown area, due to its significant correlation between context information and traffic flow [13]. To investigate the performance of CKG-GNN in different scenarios, we construct three traffic speed datasets, i.e., (1) $V_{raw}$, 06:00-10:00 for three-week workdays from 07/03/2022 to 25/03/2022, emphasizing the workday traffic pattern in morning peak hours; (2) $V_{work}$, 00:00-23:59 on Wednesday, 23/03/2022, a workday dataset; (3) $V_{rest}$, 00:00-23:59 on Saturday, 19/03/2022, a weekend dataset. The traffic speed datasets were collected every 2 minutes from HERE technologies. To enhance data quality and reduce noise, we aggregated the speed datasets into 10-min intervals by averaging. A demo of raw speed data in CCR is shown in Figure 2.
@@ IV-B Spatial KG embedding evaluation
Embedding $\mathbb{K}_{S}$ involves estimating MR’s performance with different KGE models and parameters to select the optimal configuration by capturing the complex interactions of spatial factors. In Table I, distance-based methods generally excel in embedding the spatial unit $\mathbb{K}_{S}$, as indicated by their lower both-side MR values compared to similarity-based methods. However, ComplEx, a similarity-based method, defies this trend by achieving the lowest MR (12.46 for Buffer[10-100] and Link[6]), suggesting its effectiveness in capturing spatial relationships in $\mathbb{K}_{S}$. Exclusively under the conditions of Buffer[10-100] and Link[-], TransE attains the best MR of 50.64. The superior performance of ComplEx is attributed to its ability to model asymmetric relationships through complex-valued embeddings in $\mathbb{K}_{S}$ [19]. Refer to Section V-A for the asymmetric phenomenon of the left-side and right-side evaluation. Also, the result aligns with the notion that similarity-based methods like RESCAL and NTN require a substantial amount of facts for training to perform optimally.
@@
TABLE II: Both-side MR evaluation of various spatial links in the KG embedding of $\mathbb{K}_{S}$. Link[-] is no link, and Link[1,…,6] is [Order]=1,…,6.
Model Buffer Link[-] Link[1] Link[2] Link[3] Link[4] Link[5] Link[6] ComplEx [10-100] 53.66 49.38 41.72 33.26 23.21 17.09 12.46 TransE 50.64 48.10 43.70 37.94 30.81 24.44 20.40 ComplEx [100-500] 66.52 63.36 58.60 50.69 41.04 31.90 24.86 TransE 72.89 69.90 65.62 58.34 49.07 40.36 33.25 ComplEx [10-500] 57.99 57.26 53.97 48.52 42.19 34.78 27.88 TransE 70.27 68.58 64.82 60.65 55.01 47.47 39.44
Given that both ComplEx and TransE demonstrate the potential to achieve the lowest MR in some scenarios in Table I, we further explore their performance across diverse spatial link configurations. As shown in Table II, While TransE outperforms in the specific scenario of Link[-] and Link[1] for Buffer[10-100], ComplEx demonstrates superior performance across all other configurations. A closer examination of the impact of different spatial links reveals a consistent trend: an enhancement in spatial links, as indicated by lower both-side MR values, correlates with improved embedding performance. This trend suggests that denser connections within the transportation network contribute positively to embedding $\mathbb{K}_{S}$, a pattern that holds across all buffer distances. Consequently, the optimal CKG configuration for embedding $\mathbb{K}_{S}$ appears to be Link[6] with Buffer[10-100], where ComplEx achieves the lowest both-side MR, specifically 12.46, implying its potential in providing the most suitable embeddings for subsequent traffic forecasting tasks.
@@ IV-C Temporal KG embedding evaluation
Similar to embedding $\mathbb{K}_{S}$, selecting appropriate methods and parameters to embed the temporal unit $\mathbb{K}_{T}$ is also crucial. Table III demonstrates the both-side MR results for embedding $\mathbb{K}_{T}$ using both distance-based and similarity-based methods. Notably, KG2E consistently outperforms other methods across different configurations with the lowest MR values, benefiting from its approach of representing entities and relations through multi-variate Gaussian distributions [27]. This feature makes KG2E incorporate uncertainties into the embedding process, capturing the inherent variability and complexity of relations in $\mathbb{K}_{T}$, especially in dynamic systems like traffic networks. Conversely, similarity-based methods struggle in sparse temporal connections in $\mathbb{K}_{T}$ because these methods rely on a denser graph structure to uncover and interpret the subtle patterns. Regarding the impact of [PastMins], a distinct pattern emerges: larger [PastMins] values correlate with lower MR values across all methods, suggesting that a broader time buffer enhances embedding effectiveness. The cutoff at Past[60] is strategically chosen to align with the hourly link, and extending the time buffer beyond this point could introduce redundant complexity. Therefore, Past[60] is determined to be the optimal choice. In addition, the above observations hold across two kinds of temporal links (Link[-] and Link[HDW]), reinforcing the findings’ consistency and reliability.
@@
To further investigate the most effective temporal link configuration, Table IV evaluates the both-side MR values for different temporal links with all [PastMins] using KG2E. It is noted that Link[HDW] consistently achieves the best performance, with the lowest both-side MR of 2.36 at Past[60]. Interestingly, the MR values for single temporal links—[hour], [day], and [week]—are identical. This uniformity suggests that when embedding $\mathbb{K}_{T}$, the temporal links alone (hourly, daily, or weekly) do not differentiate connections among road segments as they all evolve over time. However, when integrating all temporal links (Link[HDW]), $\mathbb{K}_{T}$ achieves an optimal structure for embedding. In summary, embedding $\mathbb{K}_{T}$ using KG2E is most effective with Link[HDW] and Past[60], offering the most effective CKG embedding for the subsequent traffic forecasting tasks.
@@ IV-D Performance of traffic speed prediction
To investigate the performance of CKG-GNN models in traffic forecasting, we utilized speed datasets $V_{raw}$ for experiments. The results, as shown in Table V, illustrate the evaluation metrics of various forecasting models over time horizons of 10, 60, and 120 minutes. The baseline GNN models establish benchmarks for comparison using the LibCity library, including TGCN, STGCN, and DCRNN [35], while the proposed CKG-GNN models integrate CKG with GNN using different context-aware units. Specifically, CKG-$\mathbb{K}_{S}$ employs the spatial unit, CKG-$\mathbb{K}_{T}$ incorporates the temporal unit, and CKG-$\mathbb{K}_{ST}$ combines both spatial and temporal units. Evaluation metrics, including MAE (mean absolute error) and MAPE (mean absolute percentage error), are presented as averages with standard deviations across five runs to present reliability and insights into the models’ consistency. For the experimental setup, we configured the epoch numbers at 500 with a batch size of 16. The optimization was performed using the Adam optimizer, initiated with a learning rate of 0.001. We applied a MultiStepLR scheduler to enhance the convergence behavior, reducing the learning rate by 0.5 at epochs 150, 250, 350, and 450. For the dual-view MHSA, we set the head numbers for the context view as 10 and for the sequence view as 4, taking into account the feature dimensions associated with each view. Furthermore, the dataset was divided into training, validation, and testing sets with ratios of 0.7:0.1:0.2, respectively.
TABLE V: Performance comparison of predicting traffic speed using different models in 10/60/120 minutes and its average values from 10 to 120 minutes using MAE (km/h) and MAPE (%). Each experiment was run five times to obtain its mean and standard deviation.
@@
MAE MAPE MAE MAPE MAE MAPE MAE MAPE
TGCN 4.82±0.04 20.66±0.30 5.34±0.03 24.10±0.17 5.52±0.06 25.34±0.31 5.32±0.03 23.92±0.15
STGCN 3.29±0.05 13.93±0.12 3.78±0.05 16.19±0.44 3.76±0.08 16.21±0.61 3.72±0.04 15.95±0.37
DCRNN 3.14±0.00 13.15±0.02 3.69±0.01 16.13±0.07 3.73±0.01 16.47±0.13 3.64±0.01 15.86±0.07
CKG-$\mathbb{K}_{S}$ 3.16±0.00 13.12±0.02 3.65±0.03 15.61±0.11 3.64±0.03 15.94±0.16 3.60±0.02 15.46±0.09
CKG-$\mathbb{K}_{T}$ 3.09±0.01 12.95±0.05 3.56±0.01 15.37±0.09 3.56±0.02 15.39±0.10 3.51±0.01 15.10±0.08
CKG-$\mathbb{K}_{ST}$ 3.08±0.01 12.88±0.05 3.52±0.00 15.04±0.15 3.47±0.02 14.86±0.11 3.46±0.01 14.76±0.09
@@
The results in Table V underscore the significant boost in forecasting performance achieved by integrating the CKG with GNN across various time horizons. When examining the average performance from 10 to 120 minutes, DCRNN stands out among the baseline models with an MAE of 3.64±0.01 and a MAPE of 15.86±0.07%. After introducing the spatial unit $\mathbb{K}_{S}$ to GNN, we observe a notable improvement compared to DCRNN, evidenced by reductions in MAE and MAPE. This enhancement is further amplified using the temporal unit $\mathbb{K}_{T}$, suggesting that temporal factors have a more pronounced impact on forecasting accuracy than spatial factors. This improvement is attributed to the direct impact of temporal factors on traffic forecasting due to their alignment with the dynamic nature of traffic flow. In contrast, spatial contexts contribute by delineating the intricate supply-demand relationship of human mobility over urban spaces, a complex aspect challenging to leverage effectively in traffic prediction models [3]. The best results are obtained when both $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$ are integrated, i.e., CKG-$\mathbb{K}_{ST}$ with the MAE and MAPE of 3.46±0.01 and 14.76±0.09%, respectively. This showcases the combined strength of spatial and temporal contexts in boosting traffic predictions.
@@
The individual performance of 10, 60, and 120 minutes reveals that the CKG-enhanced models consistently outperform the baselines, with a minor exception at the 10-min prediction using only $\mathbb{K}_{S}$. Although the 10-min MAE for $\mathbb{K}_{S}$ is slightly worse than DCRNN, the MAPE still shows an improvement. Furthermore, CKG-$\mathbb{K}_{ST}$ outperforms all the other models from 10-min to 120-min prediction, suggesting that the combination of spatio-temporal contexts contributes to the predictive reliability over both short and longer-term horizons. Regarding error progression from 10-min to 120-min forecasts, baseline models like DCRNN exhibit a gradual increase in MAE, but this error accumulation can be mitigated when context-aware knowledge graphs are integrated. Despite an increase in MAE from 10-min to 60-min forecasts across $\mathbb{K}_{S}$, $\mathbb{K}_{T}$, and $\mathbb{K}_{ST}$, their performance from 60 minutes to 120 minutes remains stable or even improves, such as the MAE decrease from 3.52±0.00 to 3.46±0.01 for CKG-$\mathbb{K}_{ST}$. This stability indicates the potent effect of CKG in enhancing longer-term traffic speed predictions.

### v2

@@ III-C2 Relation-dependent integration for context representations
The other component of the proposed relation-dependent integration strategy is attribute augmentation to enhance the semantic richness of each fact. This augmentation employs attribute facts associated with each fact in $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$, such as $hasPoi[Type]InBuffer[Dist]$ for $\mathbb{K}_{S}(poi)$ and $hasLand[Type]InBuffer[Dist]$ for $\mathbb{K}_{S}(land)$, infusing additional attribute information into the embeddings. For simplicity, we denote the attribute value for relations as $x$ and for entities as $y$. Consequently, the attribute-augmented embeddings for relations are $x_{1}\textbf{r}_{1},...,x_{l}\textbf{r}_{l}$, and for entities, it is $y\textbf{e}$. To ensure consistency and comparability across attributes, we apply max-min normalization to these attribute values separately, resulting in normalized values $x^{\prime}$ and $y^{\prime}$. Integrating these normalized attributes into the relation-dependent embedding yields the enriched context-aware embedding for $e$:
@@
Figure 2 outlines our relation-dependent integration strategy by illustrating multi-hop relation path integration and attribute augmentation for both distance-based and similarity-based models. In Figure 2(a), the objective is to generate the relation-dependent embedding for entity $\textbf{e}_{1}$ using a distance-based approach. To accomplish this, we identify a two-hop relation path $p_{1}$ associated with $\textbf{e}_{1}$, i.e., $p_{1}=r_{2}\rightarrow r_{1}$. The multi-hop relation path integration is then performed along $p_{1}$ using Equation 7, where $f{[\text{dist}]}(\cdot)=1$ and $f{[\text{sim}]}(\cdot)=0$ signifies the application of distance-based models. Meanwhile, attributes associated with each relation in the path, specifically $r_{1}$ and $r_{2}$, are incorporated into the final relation-dependent embedding using attribute augmentation, as per Equation 8. Figure 2(b) outlines a similar process for similarity-based models, with the only difference being the function where $f_{[\text{dist}]}(\cdot)=0$ and $f_{[\text{sim}]}(\cdot)=1$, reflecting the use of similarity-based models.
Figure 2: Diagrams illustrating multi-hop relation path integration and attribute augmentation for (a) distance-based models and (b) similarity-based models. The primary distinction between them lies in the method of relation path integration, with addition ($+$) employed in (a) and multiplication ($\times$) utilized in (b).
@@ III-D Context-aware traffic forecast using multi-head self-attention
In this study, we propose a novel method to apply MHSA from two views, i.e., context-based and sequence-based views, enabling a dual perspective analysis. Initially, the dual-view MHSA discerns the influence of various context features in $\mathbb{K}$, prioritizing those critical for forecasting from the context-based feature view. Subsequently, the sequence-based view employs the output from the context-based view to highlight the historical sequences vital for temporal prediction, emphasizing time slots that are most predictive for future traffic conditions. In addition, forward masking is employed in the sequence-based view to ensure temporal coherence by restricting attention to past and present data. This operation prevents the model from accessing future information, which is essential to avoid lookahead bias. In contrast to feature concatenation and traditional attention mechanisms, the dual-view MHSA method facilitates distinct and targeted processing of both contextual and sequential features, with each stream fine-tuned through dedicated attention weights. By integrating these dual perspectives, the method can synthesize insights from both views to foster a comprehensive understanding of traffic dynamics. This capability is critical for prediction tasks in complex environments where both contextual influences and historical patterns significantly impact outcomes.
Upon deriving the fused context-aware representations from the dual-view MHSA, they are subsequently integrated into a graph neural network to facilitate traffic speed forecasting, i.e., CKG-GNN. Specifically, we leverage traffic speed data alongside the context-aware representations from the preceding $a$ time steps to forecast traffic speed for the forthcoming $b$ time steps. In this study, we employ the diffusion convolution recurrent neural network (DCRNN) as our backbone GNN model, which models traffic forecasting as a diffusion process on a graph to capture both spatial and temporal dependencies through a sequence-to-sequence architecture, as detailed by [26]. When integrating the CKG framework with DCRNN, the inputs comprise two components: raw traffic speed data from the past $a$ time steps, denoted as $(v_{t-a+1},\dots,v_{t-1},v_{t})$, and context-aware representations derived from the CKG for the same time window, denoted as $(\textbf{e}^{\prime}_{fuse(t-a+1)},\dots,\textbf{e}^{\prime}_{fuse(t-1)},\textbf{e}^{\prime}_{fuse(t)})$. The output is the predicted traffic speeds for the next $b$ time steps, denoted as $(v^{\prime}_{t+1},\dots,v^{\prime}_{t+b})$. During training, the objective is to minimize the mean absolute error between the predicted traffic speed $(v^{\prime}_{t+1},\dots,v^{\prime}_{t+b})$ and actual traffic speeds $(v_{t+1},\dots,v_{t+b})$ over future time steps, i.e.,
$\min\left(\frac{1}{b}\sum_{i=1}^{b}\left|v^{\prime}_{t+i}-v_{t+i}\right|\right).$ (14)
The training is performed over 500 epochs with a batch size of 16, with the dataset divided into training, validation, and testing sets at ratios of 70%, 10%, and 20%, respectively. The optimization is conducted using an Adam optimizer with an initial learning rate of 0.001, which is adjusted by a MultiStepLR scheduler that reduces the learning rate by 50% at epochs 150, 250, 350, and 450 to enhance convergence. For the dual-view MHSA, we set the head numbers for the context view as 10 and for the sequence view as 4, tailored to the feature dimensions of each view. Furthermore, the design of our CKG-GNN model supports real-time updates to spatial and temporal contexts by exclusively changing their attributes, without altering the CKG’s structure and the embedding process. These attribute changes are directly incorporated into the inputs of $(\textbf{e}^{\prime}_{fuse(t-a+1)},\dots,\textbf{e}^{\prime}_{fuse(t-1)},\textbf{e}^{\prime}_{fuse(t)})$, which in turn impact the predicted traffic speeds.
@@ IV-A Datasets and processing
The datasets used in this study encompass three categories: traffic speed data, spatial context data, and temporal context data. All datasets used in this study are publicly accessible. Detailed information about each dataset, including their respective time frames, resolutions, and data sources, is summarized in Table A2 in the Appendix.
We sourced the traffic speed dataset from HERE technologies, focusing on Singapore’s Core Central Region (CCR), i.e., the downtown area, due to its significant correlation between contextual information and traffic flow [8]. To investigate the performance of CKG-GNN in different scenarios, we construct three traffic speed datasets, i.e., (1) $V_{raw}$, 06:00-12:00 for about three-week workdays from 08/03/2022 to 25/03/2022, emphasizing the workday traffic patterns that specifically cover the morning peak hours; (2) $V_{work}$, 00:00-23:59 on Wednesday, 23/03/2022, a workday dataset; (3) $V_{rest}$, 00:00-23:59 on Saturday, 19/03/2022, a weekend dataset. The traffic speed datasets were collected every 2 minutes from HERE technologies. To enhance data quality and reduce noise, we aggregated the speed datasets into 10-min intervals by averaging. A demo of raw speed data in CCR is shown in Figure 3.
@@ IV-B Spatial KG embedding evaluation
Embedding $\mathbb{K}_{S}$ involves estimating MR’s performance with different KGE models and parameters to select the optimal configuration by capturing the complex interactions of spatial factors. In Table I, distance-based methods generally excel in embedding the spatial unit $\mathbb{K}_{S}$, as indicated by their lower both-side MR values compared to similarity-based methods. However, ComplEx, a similarity-based method, defies this trend by achieving the lowest MR (12.46 for Buffer[10-100] and Link[6]), suggesting its effectiveness in capturing spatial relationships in $\mathbb{K}_{S}$. Exclusively under the conditions of Buffer[10-100] and Link[-], TransE attains the best MR of 50.64. The superior performance of ComplEx is attributed to its ability to model asymmetric relationships through complex-valued embeddings in $\mathbb{K}_{S}$ [16]. See Table A3 in the Appendix for the evaluation results of MRR, H@5, and AH@K, which also demonstrate the best performance of ComplEx with Buffer[10-100] and Link[6]. Also, the result aligns with the notion that similarity-based methods like RESCAL and NTN require a substantial amount of facts for training to perform optimally.
@@
TABLE II: Both-side MR evaluation of various spatial links in the KG embedding of $\mathbb{K}_{S}$. Link[-] is no spatial link, and Link[1,…,6] refers to spatial links with [Order]=1,…,6.
Model Buffer [-] [1] [2] [3] [4] [5] [6]
ComplEx [10-100] 53.66 49.38 41.72 33.26 23.21 17.09 12.46
TransE 50.64 48.10 43.70 37.94 30.81 24.44 20.40
ComplEx [100-500] 66.52 63.36 58.60 50.69 41.04 31.90 24.86
TransE 72.89 69.90 65.62 58.34 49.07 40.36 33.25
ComplEx [10-500] 57.99 57.26 53.97 48.52 42.19 34.78 27.88
TransE 70.27 68.58 64.82 60.65 55.01 47.47 39.44
Given that both ComplEx and TransE demonstrate the potential to achieve the lowest MR in some scenarios in Table I, we further explore their performance across diverse spatial link configurations. As shown in Table II, While TransE outperforms in the specific scenario of Link[-] and Link[1] for Buffer[10-100], ComplEx demonstrates superior performance across all other configurations. A closer examination of the impact of different spatial links reveals a consistent trend: an enhancement in spatial links, as indicated by lower both-side MR values, correlates with improved embedding performance. This trend suggests that denser connections within the transportation network contribute positively to embedding $\mathbb{K}_{S}$, a pattern that holds across all buffer distances. Table A4 in the Appendix also corroborates the observation that ComplEx demonstrates superior performance across most configurations, with improved outcomes when additional spatial links are incorporated, as evidenced by the metrics of MRR, H@5, and AH@K. Consequently, the optimal CKG configuration for embedding $\mathbb{K}_{S}$ appears to be Link[6] with Buffer[10-100], where ComplEx achieves the lowest both-side MR, specifically 12.46, implying its potential in providing the most suitable embeddings for subsequent traffic forecasting tasks.
@@ IV-C Temporal KG embedding evaluation
Similar to embedding $\mathbb{K}_{S}$, selecting appropriate methods and parameters to embed the temporal unit $\mathbb{K}_{T}$ is also crucial. Table III demonstrates the both-side MR results for embedding $\mathbb{K}_{T}$ using both distance-based and similarity-based methods. Notably, KG2E consistently outperforms other methods across different configurations with the lowest MR values, benefiting from its approach of representing entities and relations through multi-variate Gaussian distributions [23]. This feature makes KG2E incorporate uncertainties into the embedding process, capturing the inherent variability and complexity of relations in $\mathbb{K}_{T}$, especially in dynamic systems like traffic networks. Table A5 in the Appendix presents additional evaluation metrics: MRR, H@5, and AH@K, which suggest the selection of KG2E to ensure reliable and robust performance. Regarding the impact of [PastMins], a distinct pattern emerges: larger [PastMins] values correlate with lower MR values across all methods, suggesting that a broader time buffer enhances embedding effectiveness. The cutoff at Past[60] is strategically chosen to align with the hourly link, and extending the time buffer beyond this point could introduce redundant complexity. Therefore, Past[60] is determined to be the optimal choice. In addition, the above observations hold across two kinds of temporal links (Link[-] and Link[HDW]), reinforcing the findings’ consistency and reliability.
@@
To further investigate the most effective temporal link configuration, Table IV evaluates the both-side MR values for different temporal links with all [PastMins] using KG2E. It is noted that Link[HDW] consistently achieves the best performance, with the lowest both-side MR of 2.36 at Past[60]. Interestingly, the MR values for single temporal links—[hour], [day], and [week]—are identical. This uniformity suggests that when embedding $\mathbb{K}_{T}$, the temporal links alone (hourly, daily, or weekly) do not differentiate connections among road segments as they all evolve over time. However, when integrating all temporal links (Link[HDW]), $\mathbb{K}_{T}$ achieves an optimal structure for embedding. Refer to Table A6 in the Appendix for additional evaluation metric results of MRR, H@5, and AH@K, which also highlight the optimal performance of the Link[HDW] and Past[60] configuration for embedding $\mathbb{K}_{T}$. In summary, embedding $\mathbb{K}_{T}$ using KG2E is most effective with Link[HDW] and Past[60], offering the most effective CKG embedding for the subsequent traffic forecasting tasks.
@@ IV-D Performance of traffic speed prediction
To investigate the performance of CKG-GNN models in traffic forecasting, we utilized speed datasets $V_{raw}$ for experiments. The results, as shown in Table V, illustrate the evaluation metrics of various forecasting models over time horizons of 10, 60, and 120 minutes. We evaluate the performance of CKG-GNN models from three perspectives. First, we explore the impact of different CKG units on traffic prediction by developing three CKG-GNN variants, each integrating distinct context datasets: CKG-$\mathbb{K}_{S}$ (spatial only), CKG-$\mathbb{K}_{T}$ (temporal only), and CKG-$\mathbb{K}_{ST}$ (spatial and temporal). Second, to benchmark CKG-GNN models against baseline GNN models, we utilize TGCN, STGCN, and DCRNN from the LibCity library to predict traffic speeds over identical time horizons [32]. Third, for a stepwise comparison with other context-aware models using these context datasets, we investigate the performance of a multimodal context-based graph convolutional neural network (MCGCN) across three variants [2]: MCGCNS (spatial only), MCGCNT (temporal only), and MCGCNST (spatial and temporal). Each MCGCN model employs a DCRNN backbone with attention dimensions of 96 for both spatial and temporal contexts. Each model is run five times to ensure a robust comparison, and the results are reported as the average with standard deviations (ave±std). The employed evaluation metrics include MAE (mean absolute error), quantifying the average magnitude of prediction errors; MAPE (mean absolute percentage error), which expresses these errors as a percentage, providing a scale-independent measure of accuracy but being sensitive to small denominators; and RMSE (root mean square error), which emphasizes larger errors by squaring the prediction differences before averaging, thereby providing sensitivity to larger error magnitudes.
TABLE V: Performance comparison of traffic speed prediction across models at 10, 60, and 120 minutes, and averaged over 10 to 120 minutes, using MAE (km/h), MAPE (%), and RMSE (km/h). Each experiment was run five times to obtain its mean and standard deviation.
@@
MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE
TGCN 4.82±0.04 20.66±0.30 6.41±0.05 5.34±0.03 24.10±0.17 7.21±0.04 5.52±0.06 25.34±0.31 7.44±0.06 5.32±0.03 23.92±0.15 7.16±0.04
STGCN 3.29±0.05 13.93±0.12 4.73±0.09 3.78±0.05 16.19±0.44 5.48±0.05 3.76±0.08 16.21±0.61 5.42±0.07 3.72±0.04 15.95±0.37 5.38±0.03
DCRNN 3.14±0.00 13.15±0.02 4.50±0.01 3.69±0.01 16.13±0.07 5.37±0.01 3.73±0.01 16.47±0.13 5.40±0.01 3.64±0.01 15.86±0.07 5.27±0.01
MCGCNS 3.18±0.06 13.19±0.14 4.49±0.03 3.62±0.05 15.77±0.23 5.29±0.05 3.66±0.07 16.10±0.36 5.34±0.08 3.58±0.06 15.54±0.25 5.20±0.05
MCGCNT 3.20±0.10 13.29±0.15 4.50±0.08 3.58±0.03 15.42±0.10 5.23±0.03 3.60±0.01 15.58±0.15 5.26±0.03 3.55±0.03 15.24±0.06 5.15±0.02
MCGCNST 3.17±0.07 13.22±0.07 4.47±0.04 3.54±0.03 15.29±0.08 5.19±0.04 3.57±0.03 15.45±0.09 5.22±0.04 3.51±0.04 15.10±0.08 5.11±0.03
CKG-$\mathbb{K}_{S}$ 3.16±0.00 13.12±0.02 4.63±0.01 3.65±0.03 15.61±0.11 5.44±0.05 3.64±0.03 15.94±0.16 5.40±0.04 3.60±0.02 15.46±0.09 5.35±0.03
CKG-$\mathbb{K}_{T}$ 3.09±0.01 12.95±0.05 4.42±0.00 3.56±0.01 15.37±0.09 5.23±0.02 3.56±0.02 15.39±0.10 5.20±0.02 3.51±0.01 15.10±0.08 5.13±0.01
CKG-$\mathbb{K}_{ST}$ 3.08±0.01 12.88±0.05 4.42±0.01 3.52±0.00 15.04±0.15 5.19±0.01 3.47±0.02 14.86±0.11 5.10±0.02 3.46±0.01 14.76±0.09 5.08±0.01
The results in Table V underscore the significant boost in forecasting performance achieved by integrating the CKG with GNN across various time horizons. When examining the average performance from 10 to 120 minutes, DCRNN stands out among the baseline models with an MAE of 3.64±0.01, a MAPE of 15.86±0.07%, and an RMSE of 5.27±0.01. Using DCRNN as the backbone, both context-aware models, i.e., MCGCN and CKG-GNN, demonstrate enhanced performance when incorporating contextual information into traffic speed predictions. However, CKG-GNN consistently outperforms MCGCN when both spatial and temporal contexts are integrated. In detail, CKG-$\mathbb{K}_{ST}$ records an MAE of 3.46±0.01, a MAPE of 14.76±0.09%, and an RMSE of 5.08±0.01, marking improvements over DCRNN of 0.18 for MAE, 1.10% for MAPE, and 0.19 for RMSE. Compared to MCGCNST, the improvements are 0.05 for MAE, 0.34% for MAPE, and 0.03 for RMSE. These results demonstrate the superiority of CKG-GNN models, illustrating that they not only surpass baseline GNN models in performance but also excel beyond other advanced context-aware models. A step-by-step comparison between MCGCN and CKG-GNN reveals that MCGCNST performs suboptimally when both contexts are incorporated; however, MCGCNS, which considers only the spatial context, exhibits better MAE and RMSE than CKG-$\mathbb{K}_{S}$, although CKG-$\mathbb{K}_{S}$ achieves a better MAPE. This suggests that MCGCN’s hierarchical learning effectively captures the spatial dependencies inherent in the context. In the configuration that considers only the temporal context, CKG-$\mathbb{K}_{T}$ shows an advantage over MCGCNS, improving MAE, MAPE, and RMSE by 0.04, 0.14%, and 0.02, respectively. Overall, these findings affirm that using knowledge graphs to structure context datasets enhances the capture of spatial and temporal dependencies, thereby improving traffic prediction accuracy. Different from existing context-aware traffic forecasting methods such as MCGCN, which necessitate distinct models to accommodate the characteristics of spatial and temporal contexts [2], the proposed CKG-GNN model ensures uniformity by utilizing a consistent knowledge graph format to model both context types. Moreover, this approach not only streamlines the integration of additional context datasets without the need to develop specialized models for learning their representations but also demonstrates superior performance.
@@
Moreover, it is observed that integrating the temporal context into DCRNN generally yields greater improvements than incorporating the spatial context. This is evidenced by CKG-$\mathbb{K}_{T}$ outperforming CKG-$\mathbb{K}_{S}$ and MCGCNT surpassing MCGCNS, suggesting that temporal factors have a more pronounced impact on forecasting accuracy than spatial factors. This improvement is attributed to the direct impact of temporal factors on traffic forecasting due to their alignment with the dynamic nature of traffic flow. In contrast, spatial contexts contribute by delineating the intricate supply-demand relationship of human mobility over urban spaces, a complex aspect challenging to leverage effectively in traffic prediction models [2].
@@
The individual performance of 10, 60, and 120 minutes reveals that the CKG-enhanced models consistently outperform the baselines, with a minor exception at the 10-min prediction using only $\mathbb{K}_{S}$. Although the 10-min MAE and RMSE for $\mathbb{K}_{S}$ is slightly worse than DCRNN, the MAPE still shows an improvement. Furthermore, CKG-$\mathbb{K}_{ST}$ outperforms all the other models from 10-min to 120-min prediction, suggesting that the combination of spatio-temporal contexts contributes to the predictive reliability over both short and longer-term horizons. Regarding error progression from 10-min to 120-min forecasts, baseline models like DCRNN exhibit a gradual increase in MAE, but this error accumulation can be mitigated when context-aware knowledge graphs are integrated. Despite an increase in MAE from 10-min to 60-min forecasts across $\mathbb{K}_{S}$, $\mathbb{K}_{T}$, and $\mathbb{K}_{ST}$, their performance from 60 minutes to 120 minutes remains stable or even improves, such as the MAE decrease from 3.52±0.00 to 3.46±0.01 for CKG-$\mathbb{K}_{ST}$. This stability indicates the potent effect of CKG in enhancing longer-term traffic speed predictions.